# Generative User-Experience Research for Developing Domain-specific Natural Language Processing Applications

### v4

@@ 4.3.2 UX evaluation
User interviews
@@
List view
Search functionality
Before prototype
- The search is not efficient enough - Got used to the list view (x2) - Need to perform comparison of multiple entries and/or specific fields (x2) - Don’t know any other way - The fastest solution
(Keyword search) - Use a search to filter out a lot of non- relevant entries - A habit of googling - Know terms for searching - Determine redundancy in the records - Find solutions
After prototype
- Search for specific information, e.g., data of old analysis - Faster for some tasks
(Semantic search) - Faster and easier, especially when looking for the older entries - Easy to use - Look for specific solutions, problems, or completed tasks - Find the results faster (x2)
@@ 5 Discussion
The paper reports a case study for which we devised and implemented a methodology of making generative UX research a part of a pipeline for prototyping NLP applications. The case study showed that learning directly from users or target audiences enabled the identification of their problems and needs, which consequently led to effective and efficient prototyping with positive user feedback in the evaluation phase. Therefore, instead of evaluating hypothetical ideas for an NLP application, the concept generation phase addressed these specific needs in a problem-solving way.
We expect the proposed methodology of combining UX and NLP methodologies to impact domain-specific NLP application development positively. Understanding and evaluating data quality and availability must come in a tight combination with understanding the domain users and their stories. A timeline depicted in Figure 1 can serve as a step-by-step guideline for the generative-driven prototyping of NLP domain applications. The pipeline describes specific tasks for UX researchers, data scientists, and project managers at each stage. Combining the steps in each type of UX research and the questions from the user studies (Section 4.1 and Section 4.3) expands the guideline and equips the UX practitioners with applied examples. Section 4.1.1, Section 4.1.2, and Section 4.2 exemplify data scientists how we performed exploratory data analysis, assessed and solved the challenges for a domain NLP prototype application.

### v5

@@ 4.3.1 User-driven prototype evaluation
List view Search functionality
Before prototype - The search is not efficient enough - Got used to the list view (x2) - Need to perform comparison of multiple entries and/or specific fields (x2) - Don’t know any other way - The fastest solution (Keyword search) - Use a search to filter out a lot of non- relevant entries - A habit of googling - Know terms for searching - Determine redundancy in the records - Find solutions
After prototype - Search for specific information, e.g., data of old analysis - Faster for some tasks (Semantic search) - Faster and easier, especially when looking for the older entries - Easy to use - Look for specific solutions, problems, or completed tasks - Find the results faster (x2)
@@
User interviews: Control point and reflection
@@
4.3.2 Accuracy evaluation
The accuracy evaluation focused on the precision-oriented IR evaluation and estimated the relevance of the $N$ top-ranked results. Although such “shallow” evaluation could be incomplete, it is consistent if users are interested only in top-ranked documents in the domain application [40]. Therefore, the relevance assessment enabled the estimation of the prototype’s accuracy for daily use.
Collecting a test dataset
Controlled task execution
Document relevance level # queries $\kappa$
term level 44 .856
phrase level 16 .673
Computing accuracy metrics
After implementing the prototype, we provided two baselines to evaluate how user utility changed: a keyword search (i.e., exact term match) and BM25. The keyword search is a default SQL-based search from the logging system with which the domain users are familiar. BM25 represents a standard baseline used in IR. We implemented a more flexible version of the keyword search for the experiments. A document was retrieved if at least one term from a query was found. The documents were ranked by the decreasing number of overlapping terms and then by a timestamp of the records. For BM25, we normalized German text records and accommodated local domain setup with lemmatizing and removing stopwords. We retrieved documents if cosine similarity was higher than 0.15 and ranked them using it and a document’s timestamp. We implemented an optional context expansion in both baselines to evaluate its impact on handling the domain-specific language (see Section 4.2.2).
We evaluated the accuracy of the prototype at top $N$ records with precision-oriented metrics, i.e., precision $P$@N, and metrics for ranking quality, i.e., mean average precision $mAP$@N [40], mean reciprocal rank $MRR$ [48] and normalized discounted cumulative gain $nDCG$@N [35]. The metrics were calculated for the top 5 and 20 records in our experiments. The combined relevance scores in the test dataset were used to compute $nDCG$@N.
keyword search - 3300 .153 .068 .125 .076 .063 .112 .096
+ 5944 .132 .180 .077 .152 .101 .066 .100
BM25 - 736 .582 .294 .548 .322 .178 .479 .291
+ 657 .572 .302 .523 .337 .182 .459 .291
prototype: semantic search + 90 .888 .749 .880 .725 .668 .805 .772
Using context expansion improved precision $P$@5 and $P$@20 of both baselines. Figure 15 depicts the influence of context expansion. The ranking values measured by $MRR$, $mAP$@5, and $mAP$@20 that ignore the annotated relevance score of the documents have slightly dropped. On the contrary, $nDCG$@5 and $nDCG$@20 have increased. The results show that context expansion assisted in retrieving more candidate documents and ranking relevant documents higher.
@@ 5 Discussion
The paper reports a case study for which we devised and implemented a methodology of making generative UX research a part of a pipeline for prototyping NLP applications. The case study showed that learning directly from users or target audiences enabled the identification of their problems and needs, which consequently led to effective and efficient prototyping with positive user feedback in the evaluation phase. Therefore, instead of evaluating hypothetical ideas for an NLP application, the concept generation phase addressed these specific needs in a problem-solving way. Our case study took approximately 30 weeks to perform a full cycle to develop and collect extensive feedback on the MVP (4 weeks for exploration, 16 weeks for prototyping, and 10 weeks for evaluation). This period is rather short in terms of full-scale project development but yields better products tailored to the needs of domain experts.
We expect the proposed methodology of combining UX and NLP methodologies to positively impact domain-specific NLP application development. Understanding and evaluating data quality and availability must come in a tight combination with understanding the domain users and their stories. A timeline depicted in Figure 1 can serve as a step-by-step guideline for the generative-driven prototyping of NLP domain applications. The pipeline describes specific tasks for UX researchers, data scientists, and project managers at each stage. Combining the steps in each type of UX research and the questions from the user studies (Section 4.1 and Section 4.3) expands the guideline and equips the UX practitioners with applied examples. Section 4.1.1, Section 4.1.2, and Section 4.2 exemplify data scientists how we performed exploratory data analysis, assessed and solved the challenges for a domain NLP prototype application.